Date:
Wednesday, November 13, 2019
Telescope:
CTIO 1.5-m
Submitter:
Roberto Aviles
Submitter Email:
Observer Support:
None
Night Assistant:
None
Problems:
At 05:30 UT Moving on HIP022424 telescope stops moving in RA. In the dome with the hand paddle: fails to move E or W. It does move fine in Declination. I can hear the RA drive running but, no motion at all. Inspected the RA drive (I sent pictures to E. Parkes and J. Briones.) It seems to me a O-ring is either loose or just broken.
So, call it a night.
Program:
Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
5.75
Time Engineering:
0.0
Time lost technical:
3.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
8.75
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2019-11-13 | 5.75 | 0.0 | 3.0 | 0.0 | 0.0 | 8.75 |
Clearly hazy.